OpenStudy (anonymous):

1.) The second of three numbers is 6 more than the first. The third number is 5 less than the second. If the first number is twice as large as the third, find the 3 numbers. 2.)It took Gibbons family 2 days to travel 925 miles to their vacation home. They traveled 75 miles more on the first day than on the sedond. Hoe many miles did they travel each day? These are my last questions for the day! Thanks to everyone who is helping me!!!!

OpenStudy (campbell_st):

let x be the 1st number 2nd is x + 6 3rd is x + 6 - 5 or x + 1 if the 1st is twice as larg as the 3rd then 1/2 the 1st is equal to the 3rd 1/2 x = x + 1 solve for x

OpenStudy (campbell_st):

2nd question let x be the distance travelled on the 1st day x + 75 = distance on 2nd day x + x+75 = 925 solve for x

OpenStudy (anonymous):

can you help me solve the first one?

OpenStudy (campbell_st):

well subtract x from both sides... what do you get..?

OpenStudy (anonymous):

1/2=1

OpenStudy (campbell_st):

nope its -1/2 x = 1 now multiply both sides by -2 what do you get

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

okay thanks!!